OpenStudy (anonymous):

If , 3x^2 - 1=0 which of the following is a solution for x? ( : Can anyone show me how to solve this?

OpenStudy (anonymous):

Move the -1 to the right (it becomes positive) 3x^2=1 Then divide by 3 x^2=1/3 Now square root both sides to get rid of the square x=sqrt(1/3) or it also can be written as \[x = \frac{ \sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

Divide by 3 on both sides*